Find a) the mass in kilograms and b) rms speed for sulfur dioxide S02 at room temperature 20°C. mass5 = 32.065u


1
Expert's answer
2021-09-22T08:15:03-0400

Mass=32.065amu

"1amu=1.67\times10^{-27}kg"


"M=32.065\times1.67\times10^{-27}=5.35\times10^{-26}kg"

"V_{RMS}=\sqrt{\frac{3KT}{M}}"

"M_{so_2}=64.1\times10^{-3}kg/mol"


"V_{RMS}=\sqrt{\frac{3\times8.314\times293}{64.1\times10^{-3}}}=337.65m/sec"